# Discrete Spatial Diffusion

### v1

@@ 4.3 Lithium-ion electrodes
Electrodes in lithium-ion batteries are porous materials with a complex microstructure that governs key properties like ion transport and electrochemical performance. Nickel-manganese-cobalt cathodes, among the most common, are composed of three phases: the active material driving the electrochemical reaction, the carbon binder ensuring electrical conductivity and mechanical stability, and the pore space filled with electrolytes. The active material is expensive, creating a strong economic incentive to understand how its volume fraction and distribution influence electrode behavior. While tomographies are needed for studying microstructures and enabling computational modeling, acquiring diverse datasets is challenging (Deng et al., 2021). To overcome this, researchers often rely on computational methods to generate synthetic microstructures (Duquesnoy et al., 2023). While generative adversarial networks have been explored for this purpose, they did not control phase volume ratio parameters (Gayon-Lombardo et al., 2020). We trained a DSD model on tomography data (Usseglio-Viretta et al., 2018), where two color channels were used to represent the carbon binder and active materials. The results, shown in Fig. 5, demonstrate DSD enables precise tuning of phase volume fractions, making a powerful tool for systematically studying and optimizing electrode microstructures. For more details on datasets and reconstruction metrics applied to these samples, see Appendix G.
@@ 5 Limitations
The computational cost of forward sampling during training and reverse-time sampling during inference in DSD scales linearly with the total intensity of the image. While this makes DSD highly efficient for low-bit-depth or binary datasets, it may become less efficient than other techniques for higher-resolution images or datasets with higher intensity saturation, such as standard uint8 images. Additionally, enforcing strict intensity conservation requires a custom forward process code (Eq. (1)) and a novel sampling scheme, deviating from conventional Gaussian diffusion models. This introduces a steeper learning curve for practitioners accustomed to standard diffusion approaches. However, we argue that these trade-offs are necessary to achieve exact constraint enforcement, which is not possible with existing methods.
@@ 6 Conclusion
We introduced Discrete Spatial Diffusion (DSD), a fully discrete, mass-preserving generative model approach for images and scientific data. The foundation is to use discrete-state, continuous time statistical processes incorporating jump dynamics, rather than SDEs, as a foundation, and in particular is the first discrete diffusion model to explore spatially correlated noisification. DSD demonstrates competitive quality on standard benchmarks while enabling exact global constraints in total intensity (particle count, or mass) that are critical in many scientific applications. By preserving these constraints in both forward and reverse processes, DSD provides for exactly constrained data generation, which we explored on image synthesis and domain-specific datasets. It also demonstrates that more complex statistical processes (in this case, random walks) can be used for diffusion modeling, perhaps opening the door for further models to exploit structure in their dynamics such as conservation laws and symmetries.
Impact Statement
This paper presents work whose goal is to advance the field of generative modeling for spatial data. There are many potential societal consequences for generative modeling research, however, these are largely unspecific to the research presented here. This paper advances the ability to generate image data under constraints, which could improve the capabilities of generative models, and in particular in settings which are treated mathematically, such as physical simulations; we believe this does not broaden the scope of ethical concerns associated with generative models.
@@ Appendix A Deriving reverse-time transition rates
according to the general theory of reverse-time dynamics for continuous-time Markov systems (Campbell et al., 2022; Santos et al., 2023). We now perform the survival analysis for the inhomogeneous process. Within time ${\rm d}t$, the probability that particle $i$ leaves $(x,y,c)$ and moves to $(x+{\bar{\nu}}_{x},y+{\bar{\nu}}_{y},c)$ is $\bar{r}_{\bar{\nu}}^{i}\left(t\right){\rm d}t+\mathcal{O}\left({\rm d}t^{2}\right)$. As such, the probability of the particle remains at $(x,y,c)$ at time $t-{\rm d}t$ is $1-\sum_{{\bar{\nu}}}\bar{r}_{\bar{\nu}}^{[i]}\left(t\right){\rm d}t+\mathcal{O}\left({\rm d}t^{2}\right)$. Thanks to the independence between the particle dynamics, the probability of all $n$ particles remaining at $(x,y,c)$ at time $t-{\rm d}t$ (recall that we are evolving the reverse-time dynamics) is $1-\sum_{i=1}^{n}\sum_{{\bar{\nu}}}\bar{r}_{\bar{\nu}}^{[i]}\left(t\right){\rm d}t+\mathcal{O}\left({\rm d}t^{2}\right)$. Then, the probability of no particle leaving at a previous time $t-\Delta t$, where $\Delta t:=N{\rm d}t$ is
@@
     else if Using process likelihood then
@@
Figure 6: Structural Similarity Index Metric between the original and corrupted MNIST and CelebA images (1,000 samples evaluated at uniformly sampled random time $k\in\left\{1\ldots 2000\right\}$) with various noise scheduler. We fixed $r=120$ for MNIST and $r=200$ for CelebA in this analysis. For ours, we use Eq. (2) with $\tau_{1}=7.5$ and $\tau_{2}=2.5$. For the polynomials, $t_{k}=(k/T)^{n}$, $n=1\ldots 7$. For the cosine schedule, we use the heuristic formula given in (Nichol & Dhariwal, 2021). We remark that although the cosine schedule has been shown to be superior in previous studies (Nichol & Dhariwal, 2021; Santos & Lin, 2023), the conclusion is based on the Ornstein–Uhlenbeck process, which is distinct from the spatial diffusion process (1).
@@ Appendix C Spectral representation of DSD with periodic boundary conditions
Let us consider the two-dimensional DSD with periodic boundary conditions, whose master equation (Weber & Frey, 2017) is
@@
Appendix D Additional MNIST experiments
@@
In our additional MNIST experiments, we explored class-conditional and in-painting generation. These experiments are particularly notable due to their interactions with the mass-preserving property of DSD. For class-conditioning, we introduced class embeddings into our model following the approach described in (Song et al., 2021b). Our model performed well at the task of class retrieval, consistently producing the desired class 8. For our mass-related experiment, we tested our model on its ability to generate all of the classes given different starting masses. Because generative models struggle to extrapolate beyond training data, our model demonstrated poor performance for certain digits on masses that were too high or too low. In response to this, we picked the ’1’ with the highest mass for our high-mass test, and the ’0’ with the lowest for our low mass test, as 1 had the lowest mass of any of the numbers, and 0 had the highest. Our model performed very well on this task, consistently producing the target class even with varying mass. See Fig. 3 (d) for results.
@@
In training our model to perform in-painting, we shrunk the size of the transition matrix and held the rest of the image static. We observed high quality generations very quickly, within only 40K training steps. For our mass-related experiment, we tested the model’s reaction to increasing mass within the in-painted region and were able to see different number generations from the same starting image (Fig. 9 ).
Figure 7: Unconditional MNIST generations
Figure 8: Conditional MNIST generations
Figure 9: MNIST in-painting with progressively added mass (unconditioned)
Appendix E Hyperparameters for experiments

### v2

@@ 4.3 Lithium-ion electrodes
Electrodes in lithium-ion batteries are porous materials with a complex microstructure that governs key properties like ion transport and electrochemical performance. Nickel-manganese-cobalt cathodes, among the most common, are composed of three phases: the active material driving the electrochemical reaction, the carbon binder ensuring electrical conductivity and mechanical stability, and the pore space filled with electrolytes.
Figure 5: Generated cathode microstructures with varying phase volume fractions. The carbon binder domain appears in black, active material particles in gray, and electrolyte-filled pore space in white.
The active material is expensive, creating a strong economic incentive to understand how its volume fraction and distribution influence electrode behavior. While tomographies are needed for studying microstructures and enabling computational modeling, acquiring diverse datasets is challenging [22]. To overcome this, researchers often rely on computational methods to generate synthetic microstructures [24]. While GANs have been explored for this purpose, they do not control phase volume ratio [29]. We trained a DSD model on tomography data [74], where two color channels were used to represent the carbon binder and active materials. The results, shown in Fig. 5, demonstrate DSD enables precise tuning of phase volume fractions. We then computed key morphological metrics characterizing the electrode structure [41], demonstrating the strong generative capability of the DSD method, which generalizes across the training set more effectively than existing approaches in the literature. For more details on datasets and reconstruction metrics applied to these samples, see Appendix G.
@@ 5 Limitations
The computational cost of forward sampling during training and reverse-time sampling during inference in DSD scales linearly with the total intensity of the image. While this makes DSD highly efficient for low-bit-depth or binary datasets, it may become less efficient than other techniques for higher-resolution images or datasets with higher intensity saturation, such as standard uint8 images. While the FLOP cost of the forward noising process is higher than in conventional Gaussian diffusion models, this is handled entirely by CPU workers during the data loading phase, and so the added cost is effectively hidden completely behind the data pipeline and does not impact GPU throughput; DSD models were negligibly slower than their Gaussian counterparts, and models were trained in this work using one A100 GPU in $<\!750$ hours per model (RGB images) and $<\!50$ hours per model (microstructures, MNIST) . Inference is slower (approximately 2$\times$ per), due to the need to sample individual particle transitions, but affordable across datasets of practical interest. Despite this cost scaling, many of the largest available microstructural datasets in subsurface modeling are on the order of $1000\times 1000$ resolution. We trained DSD on such large-scale scientific data, demonstrating that the method remains computationally feasible in this regime, also suggesting the feasibility of extending it to 3D by carrying out implementation changes. Representative results from high-resolution generation are presented in Appendix F.3. Additionally, enforcing strict intensity conservation requires a custom forward process code (Eq. (1)) and a novel sampling scheme, deviating from conventional Gaussian diffusion models. This introduces a steeper learning curve for practitioners accustomed to standard diffusion approaches. However, we argue that these trade-offs are necessary to achieve exact intensity constraints, which is not possible with existing methods.
@@ 6 Conclusion
We introduced Discrete Spatial Diffusion (DSD), a fully discrete, intensity-preserving generative model approach for images and scientific data. The foundation is the use of discrete-state, continuous time statistical processes incorporating jump dynamics, rather than SDEs, and in particular is the first discrete diffusion model to explore spatially correlated noisification. DSD demonstrates competitive quality on standard benchmarks while enabling exact global constraints on particle count (thereby conserving image intensity or mass fractions under various applications) that are critical in many scientific applications. By preserving these constraints in both forward and reverse processes, DSD enables exactly constrained data generation, which we explored on image synthesis and domain-specific datasets. It also demonstrates that more complex statistical processes (in this case, random walks) can be used for diffusion modeling, opening the door for further models to exploit structure in their dynamics such as conservation laws and symmetries.
@@ Appendix A Deriving reverse-time transition rates
according to the general theory of reverse-time dynamics for continuous-time Markov systems [11, 63]. We now perform the survival analysis for the inhomogeneous process. Within time ${\rm d}t$, the probability that particle $i$ leaves $(x,y,c)$ and moves to $(x+{\bar{\nu}}_{x},y+{\bar{\nu}}_{y},c)$ is $\bar{r}_{\bar{\nu}}^{i}\left(t\right){\rm d}t+\mathcal{O}\left({\rm d}t^{2}\right)$. As such, the probability of the particle remains at $(x,y,c)$ at time $t-{\rm d}t$ is $1-\sum_{{\bar{\nu}}}\bar{r}_{\bar{\nu}}^{[i]}\left(t\right){\rm d}t+\mathcal{O}\left({\rm d}t^{2}\right)$. Thanks to the independence between the particle dynamics, the probability of all $n$ particles remaining at $(x,y,c)$ at time $t-{\rm d}t$ (recall that we are evolving the reverse-time dynamics) is $1-\sum_{i=1}^{n}\sum_{{\bar{\nu}}}\bar{r}_{\bar{\nu}}^{[i]}\left(t\right){\rm d}t+\mathcal{O}\left({\rm d}t^{2}\right)$. Then, the probability of no particle leaving at a previous time $t-\Delta t$, where $\Delta t:=N{\rm d}t$ is
@@
     else if Using likelihood loss then
@@
Figure 6: Structural Similarity Index Metric between the original and corrupted MNIST and CelebA images (1,000 samples evaluated at uniformly sampled random time $k\in\left\{1\ldots 2000\right\}$) with various noise scheduler. We fixed $r=120$ for MNIST and $r=200$ for CelebA in this analysis. For ours, we use Eq. (2) with $\tau_{1}=7.5$ and $\tau_{2}=2.5$. For the polynomials, $t_{k}=(k/T)^{n}$, $n=1\ldots 7$. For the cosine schedule, we use the heuristic formula given in [55]. We remark that although the cosine schedule has been shown to be superior in previous studies [55, 62], the conclusion is based on the Ornstein–Uhlenbeck process, which is distinct from the spatial diffusion process (1).
@@ Appendix C Spectral representation of DSD with periodic boundary conditions
Let us consider the two-dimensional DSD with periodic boundary conditions, whose master equation [77] is
@@
Appendix D Human-centric datasets
While our primary focus is on scientific data, we include results on standard vision benchmarks to contextualize DSD’s performance to the computer vision community.
D.1 MNIST experiments
As an initial validation, we trained an unconditional DSD model on MNIST and found that it reliably captures digit structure and diversity (Fig. 7).
Figure 7: Unconditional MNIST generations.
@@ D.1 MNIST experiments
Then, in our additional MNIST experiments, we explored class-conditional and inpainting generation. These experiments are particularly notable due to their interactions with the intensity-preserving property of DSD. For class-conditioning, we introduced class embeddings into our model following the approach described in [69]. Our model performed well at the task of class retrieval, consistently producing the desired class (Fig.  8).
Figure 8: Class-conditional MNIST generations for digits 0 through 4. Each row corresponds to a specified target class.
For our intensity-related experiment, we tested our model on its ability to generate all of the classes given different starting intensity. Because generative models struggle to extrapolate beyond training data, our model demonstrated poor performance for certain digits on initial intensity values that were too high or too low. In response to this, we picked the ’1’ with the highest intensity for our high-intensity test, and the ’0’ with the lowest for our low intensity test, as 1 had the lowest intensity of any of the numbers, and 0 had the highest. Our model performed very well on this task, consistently producing the target class even with varying intensity. See Fig.  3 (d) for results.
@@
In training our model to perform inpainting, we shrunk the size of the transition matrix and held the rest of the image static. We observed high quality generations very quickly, within only 40K training steps. For our intensity-related experiment, we tested the model’s reaction to increasing intensity within the inpainted region and were able to see different number generations from the same starting image (Fig. 9 ). See section E for comparisons with other conditioning approaches.
Figure 9: Unconditional MNIST inpainting with progressively increasing total intensity in the masked region. Rows correspond to increasing intensity levels, illustrating how digit identity and structure evolve under fixed spatial context.
D.2 CIFAR-10 dataset
We trained DSD on the CIFAR-10 dataset and evaluated its generative performance under different sampling configurations. Specifically, we tested four values of the Courant–Friedrichs–Lewy (CFL) tolerance parameter, $\varepsilon\in{0.01,0.05,0.10,0.15}$, which controls the reverse-time integration step size. As expected, lower values of $\varepsilon$ yield more accurate reverse-time integration and lead to improved Fréchet Inception Distance (FID).
While FID is not the primary focus of our model, the discrete-state pixel-hoping nature of DSD introduces local high-frequency fluctuations during sampling (Fig. 10) that reduce smoothness. These artifacts, though minor, negatively impact FID and perceptual metrics. To address this, we apply a lightweight bilateral filter as a post-processing step to each generated image. This filter is non-trainable, respects both spatial proximity and intensity similarity, and acts as a smooth denoising operator that preserves semantic edges and attenuates visually incoherent fluctuations. Importantly, we apply it with default hyperparameters across all images, without per-sample tuning. The filtering operation is computationally negligible and can be implemented as a fixed convolutional layer at the end of the generation pipeline.
This simple step substantially improves both FID and spatial FID (sFID), without degrading the underlying structure of the samples (Fig.12). In Fig.11b, we quantify the impact of filtering across CFL tolerances. As an additional refinement step, we also use a pre-trained off-the-shelve CIFAR-10 classifier to discard semantically incoherent samples. Filtering based on classifier confidence can be integrated early in the sampling loop and correlates strongly with out-of-distribution or degenerate outputs (Fig. 11c).
A summary of results of these sampling experiments is shown in Table 1. While an FID score of around 20 does not represent state-of-the-art performance, we hypothesize that incorporating larger architectures, advanced optimization strategies, multi-GPU trainings, and data augmentation could further reduce this gap. DSD may be a viable foundation for discrete, constraint-preserving generative modeling even in conventional RGB settings.
Figure 10: Unconditional CIFAR-10 generations produced by DSD with CFL tolerance $\varepsilon=0.01$. Samples are generated without any post-processing. While structurally coherent, mild pixel-level irregularities are visible.
Figure 11: FID scores computed on 20,000 CIFAR-10 generations under different sampling configurations. (Left:) Samples generated with different CFL tolerances ($\varepsilon$). (Middle:) Same samples after applying a fixed bilateral filter. (Right:) Bilaterally filtered samples sampled with 1% further refined by discarding outputs with classifier confidence below a certain threshold.
Figure 12: CIFAR-10 generations after bilateral filtering. These images correspond directly to the unfiltered samples shown in Fig. 10, post-processed using a fixed, non-trainable bilateral filter. The operation smooths pixel-level irregularities introduced by the discrete sampling process while preserving structural and semantic content.
Table 1: Quantitative evaluation of CIFAR-10 generations under different sampling and filtering strategies. The CFL tolerance $\varepsilon$ controls the reverse-time integration step size. Filtering refers to post-processing with a bilateral filter.
Sampling Configuration Images FID $\downarrow$ sFID $\downarrow$
CFL $\varepsilon=0.01$ 50,000 46.3 23.4
CFL $\varepsilon=0.01$ + filtering 50,000 28.2 16.4
+ Classifier thresholding ($p\geq 0.99$) 21,780 20.6 17.7
D.3 CelebA dataset
To demonstrate scalability to higher-resolution, human-centric data, we trained DSD on the CelebA 64$\times$64 dataset. While the resulting FID of 29 is not competitive with state-of-the-art models, it remains consistent with our CIFAR-10 results and underscores the feasibility of applying DSD to RGB images with more complex structure. As illustrated in Figure 13, the model is capable of producing globally coherent facial samples, despite operating in a discrete, intensity-conserving setting. We speculate that such discrete spatial models could be beneficial for downstream applications involving controlled image manipulation, such as recolorization, region editing, or consistency-preserving transformations.
Figure 13: Unconditional generations of the CelebA 64$\times$64 dataset.
Table 2: Quantitative evaluation of CelebA 64$\times$64 generations under different sampling strategies. The CFL tolerance $\varepsilon$ controls the reverse-time integration step size. Filtering refers to post-processing with a bilateral filter. These scores were computed with 50,000 realizations.
Sampling Configuration FID $\downarrow$ sFID $\downarrow$
CFL $\varepsilon=0.05$ 44.2 29.0
CFL $\varepsilon=0.05$ + filtering 28.9 25.7
Appendix E Comparison with existing frameworks
E.1 Gaussian Diffusion with Conditioning
To benchmark against conventional generative models, we implemented a Gaussian diffusion baseline conditioned on the target total intensity (or mass). The architecture employed a continuous-time variant of NCSN++ with explicit conditioning on total intensity per image via an auxiliary embedding channel. Fig. 14 shows the behavior of this model across a range of target intensities. Near the dataset mean, where conditioning remains within the empirical distribution, the model performs moderately well. However, closer examination reveals that the generated images often contain small negative values, even though training data was strictly bounded in $[0,1]$. While this behavior may be inconsequential in floating-point representations, it becomes problematic when the outputs are discretized and clipped for evaluation or downstream use. In particular, thresholding negative values and rounding to 8-bit precision systematically increases the intensity error. These issues become increasingly severe in the distribution tails, where the model frequently fails to respect the specified conditioning.
In contrast, our proposed DSD framework guarantees exact intensity preservation by construction, without requiring post hoc corrections, which holds uniformly across the intensity distribution, including extreme quantiles, as shown in Fig. 14. Notably, when sampling at the lowest intensity levels, DSD generates extremely thin or skeletal "1"s, while very high intensities produce over-saturated, bold structures such as "fat" zeros that span much of the image domain. This behavior suggests that the model is not merely enforcing conservation, but learning meaningful structural adaptations to accommodate different global constraints, consistent with underlying patterns in the data.
Figure 14: Sample-wise mean intensity across 1,000 realizations of mnist digits for each target intensity. We compare Gaussian diffusion (white), its discretized version (gray), and Discrete Spatial Diffusion (black horizontal lines). Each violin shows the distribution of mean intensities computed across 1,000 individual samples.
Our observations indicate that: (1) intensity can be statistically conditioned within the distributional range, but without guarantee of exactness; (2) rounding and clipping of generated samples introduce further bias, revealing that the model relies on physically invalid negative intensities to approximate the constraint; and (3) extrapolation outside the data manifold results in complete failure of conditioning.
E.2 Guidance Function Approach
In reinforcement learning settings, a guidance function can be naturally defined using task-specific reward or cost structures [40]. In generative modeling, however, such functions are generally unavailable, and often must be learned or heuristic in nature, as discussed in work on posterior sampling. Specifically, exact guidance is only possible when the conditional distribution as exact posteriors (of generative samples given the constraint of interest) is known, which is not the case for total intensity constraints in image generation.
To explore whether heuristic guidance functions could serve as a surrogate, we conducted additional experiments applying a mean-squared error (MSE) penalty between the generated and target total intensity. Our results showed that while guidance can reduce error somewhat for intensities near the mean training data intensity, it fails to enforce conditioning reliably, and especially so at the tails of the training distribution. In these cases, samples frequently severely violate the intensity constraint.
Appendix F Subsurface microstructures
F.1 Detailed description of X-ray scans of subsurface rocks
  • Berea Sandstone: This sandstone sample from [54] provides a high-resolution image of the rock microstructures obtained through X-ray microtomography (X-ray $\mu$CT). In this process, the rock sample is rotated while being scanned by an X-ray beam, capturing a series of 2D radiographs at different angles. These projections are then computationally reconstructed into a 3D volume, where each voxel represents the X-ray attenuation of the material at that location. The X-ray microtomography scans were performed using a SkyScan 1272 system, operating at 50 kV and 200 $\mu$A, with a CCD detector capturing projections at a resolution of 2.25 $\mu$m per voxel. The resulting dataset consists of grayscale images with a voxel size of 2.25 $\mu$m, where variations in intensity distinguish between the solid matrix and the pore space. The solid matrix primarily consists of tightly packed mineral grains—mostly quartz—while the pores are voids that can be occupied by fluids such as water or hydrocarbons. After preprocessing steps like contrast enhancement, noise reduction, and segmentation, the final dataset represents the pore network. The Berea sample has a measured porosity of 18.96% and permeability of 121 mD. This dataset is particularly useful for computational modeling, as it enables direct comparison between numerical simulations and experimentally measured permeability, providing a rich testbed for learning-based methods that seek to map complex microstructural information to macroscopic transport properties. This sedimentary rock is a well-characterized geological benchmark, widely used in fluid flow studies due to its homogeneous grain structure and consistent permeability properties, making it a good first benchmark for our study.
  • Savonnières Carbonate: This carbonate sample, described in [9], is a layered, oolithic grainstone with a wide porosity and a permeability varying from 115 to over 2000 mD, depending on local heterogeneities. The rock is characterized by a highly multimodal and interconnected pore structure, with distinct macropores and microporosity. X-ray microtomography (X-ray $\mu$CT) was used to image the sample at a resolution of 3.8 $\mu$m voxel size, revealing intricate pore geometries. The sample was scanned at the Ghent University Centre for X-ray Tomography (UGCT) using their HECTOR scanner, developed in collaboration with XRE, Belgium. The macropores include both intergranular voids and hollow ooids, while the microporosity is found within ooid shells and intergranular spaces. Micropores in the sample often serve as the primary pathways connecting poorly connected macropores, creating a complex hierarchical network. After preprocessing steps, including noise reduction, anisotropic diffusion filtering, and watershed segmentation, a multiscale pore network model was extracted. This dataset is particularly compelling due to its extreme heterogeneity, with pore sizes spanning orders of magnitude, and its ability to represent coupled serial and parallel flow pathways. Savonnières serves as a test case for studying the impact of complex samples in our workflow.
  • Massangis Limestone: This oolitic limestone sample from [8] is a highly heterogeneous carbonate rock with a complex, multimodal pore structure resulting from diagenetic alterations, including dolomitization and dedolomitization. The rock contains a mix of intergranular and moldic macroporosity, along with microporosity concentrated in ooid rims and partially dissolved dolomite regions. Its porosity ranges from 9.5% to 13.8%, depending on local variations, and its permeability is highly anisotropic due to the interplay between connected macropores and poorly accessible microporosity. X-ray microtomography (X-ray $\mu$CT) was used to image the sample at a voxel resolution of 4.54 $\mu$m, capturing the intricate connectivity of macro- and micropores. The sample was scanned at the Ghent University Centre for X-ray Tomography (UGCT) using a FeinFocus FXE160.51 transmission tube, in collaboration with Paul Scherrer Institute (PSI), Switzerland. Differential imaging was applied to enhance the detection of fluid-filled microporosity, revealing the rock’s internal heterogeneities. Unlike more uniform carbonate samples, Massangis exhibits significant spatial variations in pore connectivity, leading to zones of high permeability interspersed with isolated pore networks. This dataset serves as another challenging benchmark for modeling porous media microstructure.
F.2 Quantitative metrics of the generated images
In porous media analysis, characterizing the spatial arrangement and size distribution of pores is crucial for understanding transport properties, mechanical behavior, and overall structure-function relationships. To quantify these characteristics, we compute the spatial correlation function and pore size distribution (PSD) using PoreSpy [34], a Python-based toolkit for quantitative analysis of porous media images. The Pore Size Distribution (PSD) characterizes the variation of pore sizes within a porous material, providing insights into connectivity, permeability, and flow dynamics. The most common method to determine PSD computationally is the local thickness approach. Given a binary image $I(x,y)$, where pore space is represented as 1 and solid space as 0, the pore size function $f(r)$ is defined as the probability density function (PDF) of the largest sphere that can be inscribed at any point within the pore space. The PSD provides a statistical summary of pore connectivity and transport properties. Small pores dominate permeability, while large pores govern bulk flow. Both of these metrics for the training and generated samples are shown in Fig. 16, our realizations exhibit excellent agreement with the training data.
Figure 16: Quantitative comparison between training and generated rock samples. (Left) Two-point correlation function, showing excellent agreement of spatial features between training data (gray) and 100 randomly generated samples (blue). (Middle) Normalized pore size distribution schematic, with colors indicating relative pore sizes. (Right) Cumulative distribution function (CDF) of the normalized pore sizes, comparing the statistical distribution of training and generated samples.
Additionally, following the evaluation protocol in Lee and Yun [46], we computed Fréchet Inception Distance (FID) for our generated carbonate samples, obtaining a score of 0.9—substantially lower than their reported value of 18.1 on rocks of similar lithology. While this comparison is not strictly one-to-one due to dataset and preprocessing differences, it offers a compelling quantitative indication of the improved realism and fidelity of DSD-generated microstructures. We note, however, that FID may have limited interpretability for binary porous media images, and should be interpreted alongside domain-specific metrics.
F.3 Large scale generations.
To evaluate the scalability of DSD in realistic scientific settings, we trained a model on high-resolution binary microstructure data of Leopard Sandstone, an interesting sample from the geological standpoint. This dataset consists of $1000\times 1000$ samples representing complex subsurface structures characterized by heterogeneity, anisotropy, and nontrivial pore connectivity. Generating such large samples is particularly relevant for studying representative elementary volumes.
Despite the increased resolution and number of particles, DSD remains computationally feasible in this regime. The forward process remains efficiently parallelizable across CPU workers, and reverse-time sampling scales linearly with intensity while maintaining tractable runtimes. A subset of generated samples is shown in Fig. 17, demonstrating the model’s ability to reproduce structurally diverse and geologically plausible microstructures at scale.
While extending DSD to 3D requires modest modifications to the implementation (e.g., defining a 6-connected lattice for particle jumps), we find the approach remains tractable for volumetric data, further supporting its suitability for scientific applications in porous media and materials modeling.
Figure 17: Realizations of Discrete Spatial Diffusion (DSD) on 1,000$\times$1,000 images of Leopard Sandstone. These samples demonstrate that our model scales to high-resolution, structurally complex porous media. Leopard Sandstone exhibits heterogeneous, anisotropic pore structures representative of real-world geological variability.
Appendix G X-ray scans of NMC cathodes
G.1 Dataset description
This dataset provides high-resolution 3D images of a Li-ion battery cathode composed of active material (nickel-manganese-cobalt oxide, NMC), carbon black, and a polymer binder [74]. The cathode sample was imaged via X-ray microtomography (X-ray µCT) and nano-tomography (X-ray nano-CT) to capture both the overall electrode architecture and fine-scale features of the carbon/binder domain (CBD). For micro-CT, a Zeiss Xradia Versa 520 system was operated at 80 kV and 88 µA, acquiring projections at an effective isotropic voxel size of approximately 398 nm over a field of view of about 400 µm. The nano-CT scans were performed using a Zeiss Xradia Ultra 810 system with a chromium target (35 kV, 25 mA), yielding isotropic voxel sizes on the order of 126 nm across a field of view of approximately 64 µm. In both cases, the 2D radiographs were reconstructed into 3D grayscale volumes using a filtered back-projection algorithm, capturing the X-ray attenuation due to the dense NMC particles and the less attenuating pore/CBD regions.
G.2 Effective metrics
We computed these metrics using the Python library TauFactor [41], and the comparisons between the real and generated images based on these metrics are illustrated in Fig. 19, while a collection of the training data and generated images is in Fig.18.
Figure 18: Microstructural characterization metrics for 80 training samples and 80 generated samples. The boxes show the 25th-50th-75th percentile, the whiskers the minimum and maximum values. Metrics computed using TauFactor [41].
Appendix H Hyperparameters for experiments